# Second Generation 3-DAP 3-Dimensional Assignment Problem

### v2

@@
All the authors are with the School of Electrical Engineering and Computer Science KTH Royal Institute of Technology, Stockholm, Sweden (e-mail: xiay@kth.se, sraz@kth.se, carlofi@kth.se). C. Fischione is also with Digital Futures of KTH.
@@ Abstract
In this paper, we consider digital over-the-air computation (AirComp) and introduce a new multi-symbol modulation framework called sequential modulation for AirComp (SeMAC). Building upon ChannelComp, a general framework for designing modulation schemes to support arbitrary function computation over a multiple access channel (MAC), SeMAC maps each input value to a sequence of modulated symbols using distinct constellation diagrams across multiple time slots. This extension generalizes ChannelComp by enabling flexible modulation design across multiple transmissions, thereby enhancing reliability against channel noise. We formulate the modulation design as a non-convex optimization problem, apply matrix lifting to relax it into a semidefinite programming (SDP), and recover a feasible modulation solution by solving a low rank approximation. For scenarios where the modulation formats cannot be changed, we further develop a power adaptation scheme that adjusts amplitude and phase of the modulated symbols while preserving the modulation structure. Numerical results show that SeMAC can achieve a reliable computation by reducing the computation error up to 18 dB compared to other existing methods, particularly for the product function.
@@ I Introduction
With the widespread deployment of Internet of Things (IoT), cellular networks are shifting from “connected everything” in 5G to “connected intelligence” in 6G systems [1]. This shift is driven by the growing need to process distributed data at the network edge devices [2]. In this context, AirComp has emerged as a promising technique that leverages the signal superposition property to compute functions directly during transmission [3]. The information-theoretic studies [4] shows that such joint computation-communication designs can outperform separating communication from computation in terms of efficiency and latency.
@@
However, analog AirComp relies on uncoded transmission, making it vulnerable to noise and fading, and incompatible with modern digital infrastructures [5]. To address these limitations, several digital AirComp methods have been developed, such as one-bit broadband digital aggregation (OBDA), majority-vote frequency-shift keying (FSK), and their asynchronous and non-coherent extensions. However, these methods are often limited to fewer number of transmitters or specific functions [3]. To provide a more general and scalable solution, ChannelComp [6] was introduced to enable arbitrary function computation over the MACs by designing digital modulation schemes via optimization. To reduce its optimization complexity, SumComp [7] focuses on summation function computation by jointly designing the input coding and constellation mapping. Both frameworks adopt the single-symbol modulation design, where each input is mapped to a single modulated symbol.
While the digital AirComp solutions make it well-suited for practical deployment, its reliability under noisy conditions is not fully guaranteed. As a result, recent works has focused on multi-symbol modulation to enhance the computation performance through redundant symbol transmission [8, 9, 10]. Bit-Slicing [8] improves robustness by dividing the bit representation of each input into segments, transmitted as separate modulated symbols. ReMAC [9] enhances reliability by selectively repeating modulated symbols across multiple time slots. However, Bit-Slicing suffers from increased approximation error when extended to nonlinear functions, while the repetition coding design in ReMAC leads to higher computational complexity [9].
In this paper, we propose SeMAC, a multi-symbol modulation framework for providing a reliable computation. The key idea is to encode each input value into a sequence of modulated symbols using distinct constellation digrams, as opposed to ReMAC which uses an identical constellation pattern across all time slots. This distinction offers greater degrees of freedom in modulation design, thereby yielding a more robust computation performance. To enable this, we formulate an optimization problem to design the multi-symbol modulation. Due to the intractability of this problem, we relax it into a SDP using lifting technique, and the final modulation design is obtained through a low rank approximation of the lifted solution. In addition, we consider scenarios where the modulation pattern is fixed. In this case, we introduce a power adaptation scheme, called SeMAC with power adaptation (SeMAC-PA), which adjusts the power and phase of the existing modulation structure for reliable computation. Numerical results show that SeMAC enables robust computation over noisy MACs and reduces computation error by up to 18 dB compared to ReMAC, particularly for the product function.
@@
Figure 1: The overall process of proposed SeMAC via multi-symbol modulation design. Given the input set $x_{1}^{(i)},\ldots,x_{K}^{(i)},i\in[M]$, each node $k$ encodes its input value $x_{k}^{(i)}$ into a modulation sequence $\mathcal{E}_{k}(x_{k}^{(i)})=[\vec{x}_{k,1}^{(i)},\ldots,\vec{x}_{k,L}^{(i)}]$. At each time slot $\ell\in[L]$, the CP receives a signal given by $\vec{y}_{\ell}^{(i)}=\sum_{k=1}^{K}h_{k,\ell}p_{k,\ell}\vec{x}_{k,\ell}^{(i)}+\vec{z}_{\ell}$, where $h_{k,\ell}$ and $p_{k,\ell}$ denote the channel coefficient and transmit power of node $k$ at time slot $\ell$, respectively, and $\vec{z}_{\ell}$ is the additive white Gaussian noise at time $\ell$. Finally, the tabular function $\mathcal{T}(\cdot)$ maps the aggregated sequence $\vec{y}_{1}^{(i)},\ldots,\vec{y}_{L}^{(i)}$ into the function output value $f^{(i)}$.
@@ II-A Signal Model
Consider a network with $K$ single-antenna nodes transmitting data to a computation point (CP) over a shared channel. Each node $k$ sends a data value $x_{k}\in\mathbb{F}_{Q}$, where $\mathbb{F}_{Q}$ denotes a finite field with $Q=2^{b}$ elements, where $x_{k}$ is represented using $b$ bits. The goal of the CP is to compute the desired function $f(x_{1},\ldots,x_{K})$ over the MACs. Each node encodes its data value $x_{k}$ into a modulated signal $\vec{x}_{k}$ through an encoder $\mathcal{E}_{k}$, such that $\vec{x}_{k}=\mathcal{E}_{k}(x_{k})\in\mathbb{C}$. We assume perfect synchronization of carrier frequency and symbol timing among all the nodes and the CP, and then the CP receives the superimposed signal as
@@
Assuming the availability of perfect channel state information, each node applies optimal power control by inverting the channel [11], given by $p_{k}=h_{k}^{*}/|h_{k}|^{2}$. In this regard, we can compensate for channel distortion and ensure coherent signal alignment at the CP. With channel inversion applied, and ignoring the effect of the noise, the received signal $\vec{r}:=\sum\nolimits_{k=1}^{K}\vec{x}_{k}$ yields a finite constellation diagram. Then, the CP applies a tabular function $\mathcal{T}(\cdot)$ to map each constellation point to its corresponding function output.
@@ II-B Overlap Avoidance
$|\vec{v}^{(i)}-\vec{v}^{(j)}|\geq\epsilon|f^{(i)}-f^{(j)}|^{2},\quad\forall(i,j)\in[M]^{2},$ (2)
where $M$ is the cardinality of the function range, and $\epsilon>0$ is a positive constant.
@@ III-A Multi-Symbol Transmission
Furthermore, we construct the modulation matrix $\bm{X}:=[\bm{X}_{1}^{\mathsf{T}},\dots,\bm{X}_{K}^{\mathsf{T}}]^{\mathsf{T}}\in\mathbb{C}^{N\times L}$, where $N:=QK$, by vertically stacking the modulation matrices of all $K$ nodes. Give the input combination $x_{1}^{(i)},\dots,x_{K}^{(i)}$, we define a binary vector $\bm{a}_{i}\in\{0,1\}^{N}$ which consists of $K$ one-hot blocks of length $Q$, each selecting the encoded modulation symbol at one node. Therefore, the resultant modulation sequence associated with the function output $f^{(i)}$ is given by $\bm{v}^{(i)}=\bm{a}_{i}\bm{X}\in\mathbb{C}^{1\times L}$.
@@ III-B Multi-Symbol Estimation and Decoding
$\|(\bm{a}_{i}-\bm{a}_{j})^{\mathsf{T}}\bm{X}\|_{2}^{2}\geq\epsilon|f^{(i)}-f^{(j)}|,\quad\forall(i,j)\in[M]^{2},$ (6)
where $\|\cdot\|_{2}$ denotes the $\ell_{2}$-norm. Note that if the condition in (6) is not satisfied, the corresponding Voronio regions are merged, and we consider the average of the associated function outputs as the final function output [6]. The overall SeMAC process is shown in Fig. 1.
@@ IV-A Modulation Design With Channel Inversion
In this subsection, the modulation scheme is designed under the assumption of channel inversion power control. Given that $Q>L$, each modulation matrix $\bm{X}_{k}$ is subject to a rank constraint as ${\rm rank}(\bm{X}_{k})\leq L$. To avoid the overlaps among the sequence of the constellation points, we then pose the following feasibility optimization problem to enable robust computation, subject to the norm power constraint:
$\mathcal{P}_{0}:=$ ${\rm find}\quad\bm{X}$
${\rm s.t.}\quad$ $\|(\bm{a}_{i}-\bm{a}_{j})^{\mathsf{T}}\bm{X}\|_{2}^{2}\geq\Delta f_{i,j},\quad\forall(i,j)\in[M]^{2},$ (7a)
${\rm rank}(\bm{X}_{k})\leq L,$ (7b)
$\|\bm{x}_{k,\ell}\|_{2}^{2}\leq 1,\quad\forall k\in[K],\quad\forall\ell\in[L],$ (7c)
where $\Delta f_{i,j}:=\epsilon|f^{(i)}-f^{(j)}|$. However, $\mathcal{P}_{0}$ is a non-convex quadratically constrained quadratic programming (QCQP), and solving it is NP-hard [6]. To address this non-convexity, we apply a lifting technique by introducing a new matrix variable $\bm{W}:=\bm{X}\bm{X}^{\mathsf{H}}\in\mathbb{C}^{N\times N}$, which is positive semidefinite and has the following block structure:
$\bm{W}=\begin{bmatrix}\bm{W}_{1}&\bm{X}_{1}\bm{X}_{2}^{H}&\dots&\bm{X}_{1}\bm{X}_{K}^{H}\\ \bm{X}_{2}\bm{X}_{1}^{H}&\bm{W}_{2}&\dots&\bm{X}_{2}\bm{X}_{K}^{H}\\ \vdots&\vdots&\ddots&\vdots\\ \bm{X}_{K}\bm{X}_{1}^{H}&\bm{X}_{K}\bm{X}_{2}^{H}&\dots&\bm{W}_{K}\end{bmatrix}$, (8)
where each diagonal block $\bm{W}_{k}:=\bm{X}_{k}\bm{X}_{k}^{\mathsf{H}}\in\mathbb{C}^{Q\times Q}$ is also positive semidefinite. This lifting step transforms the non-convexity into a rank constraint ${\rm rank}(\bm{W}_{k})\leq L$. Then, by dropping the rank constraint of each $\bm{W}_{k}$, Problem $\mathcal{P}_{0}$ is relaxed into a convex SDP as follows:
$\mathcal{P}_{0}\approx\mathcal{P}_{1}:=$ ${\rm find}\quad{\bm{W}}$
${\rm s.t.}\quad$ ${\rm Tr}(\bm{W}\cdot\bm{D}_{i,j})\geq\Delta f_{i,j},\quad\forall(i,j)\in[M]^{2},$ (9a)
$\bm{W}_{k}\succeq\bm{0},\quad{\rm Tr}(\bm{W}_{k})\leq L,\quad\forall k\in[K],$ (9b)
where $\bm{D}_{i,j}=(\bm{a}_{i}-\bm{a}_{j})(\bm{a}_{i}-\bm{a}_{j})^{\mathsf{T}}$. Problem $\mathcal{P}_{1}$ can be solved by convex solver tools such as CVX [12], yielding an optimal solution $\bm{W}^{*}$. Let $L_{k}^{*}={\rm rank}(\bm{W}_{k}^{*})$ denote the rank of each digonal block. Since $\bm{W}_{k}^{*}$ may not satisfy the removed rank constraint, a low rank matrix approximation [13] is applied to project each $\bm{W}^{*}_{k}$ back onto the feasible region of $\mathcal{P}_{0}$.
$\mathcal{P}_{2}:=$ $\arg\min_{\bm{X}}\quad\|\bm{W}^{*}_{k}-\bm{X}_{k}\bm{X}_{k}^{\mathsf{H}}\|_{\mathsf{F}}^{2}$
${\rm s.t.}\quad$ ${\rm rank}(\bm{X}_{k})\leq L,$ (10a)
$\|\bm{x}_{k,\ell}\|_{2}^{2}\leq 1,\quad\forall k\in[K],\quad\forall\ell\in[L].$ (10b)
To solve Problem $\mathcal{P}_{2}$, a singular value decomposition [14] is applied to each matrix $\bm{W}_{k}^{*}$, expressed as $\bm{W}_{k}^{*}=\bm{U}_{k}\bm{\Sigma}_{k}\bm{U}_{k}^{\mathsf{H}}$. Here, $\bm{U}_{k}\in\mathbb{C}^{Q\times Q}$ is a unitary matrix with columns $\{\bm{u}_{k,1},\ldots,\bm{u}_{k,Q}\}$ representing the singular vectors, and $\bm{\Sigma}_{k}\in\mathbb{R}^{Q\times Q}$ is a diagonal matrix containing the singular values $\{\sigma_{k,1},\ldots,\sigma_{k,Q}\}$ in descending order. Based on this decomposition, the optimal solution to Problem $\mathcal{P}_{2}$ is constructed as $\bm{X}_{k}^{*}$, which yields the modulation design for node $k$, as presented in the following proposition.
Proposition 1.
The optimal solution to optimization problem $\mathcal{P}_{2}$ for each node $k$ is given by
$\bm{X}^{*}_{k}=\sum\nolimits_{i=1}^{\min(L,L^{*}_{k})}s_{k,i}\bm{u}_{k,i}\bm{e}_{i}^{\mathsf{T}},$ (11)
where the scaling factor $s_{k,i}$ is defined as
$s_{k,i}=\min\left(\sqrt{\sigma_{k,i}},1\right),\quad\forall k\in[K].$ (12)
Here, $\bm{e}_{i}\in\{0,1\}^{N\times 1}$ is the elementary vector with a 1 in the $i$-th entry and 0 elsewhere.
Proof.
See Appendix -A. ∎
@@
Note that $\bm{X}_{k}^{*}$ serves as an approximate solution to the original modulation design problem $\mathcal{P}_{0}$, and is obtained by projecting the optimal lifted matrix $\bm{W}_{k}^{*}$ from the relaxed SDP problem $\mathcal{P}_{1}$ onto a rank-constrained space in $\mathcal{P}_{2}$. This approximation retains the dominant rank-$L$ singular components of $\bm{W}_{k}^{*}$, as shown in Proposition 1. When $\bm{W}_{k}^{*}$ is close to rank-$L$, the spectral tail $\sum_{i>L}\sqrt{\sigma_{k,i}}$ becomes negligible [15], and hence $\bm{X}_{k}^{*}$ closely satisfies the constraint (7a) in $\mathcal{P}_{0}$.

### v3

@@
FOMs
first-order methods
FPP-SCA
Feasible Point Pursuit-Successive Convex Approximation
@@
All the authors are with the School of Electrical Engineering and Computer Science KTH Royal Institute of Technology, Stockholm, Sweden (e-mail: {xiay,sraz,carlofi}@kth.se). C. Fischione is also with Digital Futures of KTH.
@@ Abstract
Recently, over-the-air computation (AirComp) leverages the superposition property of wireless channels to enable efficient function computation over a multiple access channel (MAC). However, existing digital AirComp methods either rely on single-symbol modulation, which limits flexibility and robustness, or on multi-symbol extensions that suffer from high complexity or approximation errors. To overcome these limitations, we propose a new multi-symbol modulation framework, termed sequential modulation for AirComp (SeMAC), which encodes each input into a sequence of symbols with distinct constellation diagrams across multiple time slots. This approach increases design flexibility and robustness against channel noise. Specifically, the modulation design is formulated as a non-convex optimization problem and efficiently solved through a successive convex approximation (SCA) combined with stochastic subgradient descent (SSD). For fixed modulation formats, we further develop SeMAC with power adaptation (SeMAC-PA) to adjusts transmit power and phase while preserving the modulation structure. Notably, numerical results show that SeMAC improves computation accuracy by up to 14 dB compared to the existing methods for computing nonlinear functions such as the product function.
@@ I Introduction
With the widespread deployment of Internet of Things (IoT), cellular networks are shifting from “connected everything” in 5G to “connected intelligence” in 6G systems [1]. This shift is driven by the growing need to process distributed data at the network edge devices [2]. In this context, AirComp has emerged as a promising technique that leverages the signal superposition property to compute functions directly during transmission [3]. In particular, AirComp supports 6G intelligence applications such as collaborative sensing and over-the-air federated learning (AirFL), where local model updates are efficiently aggregated across devices to enable global model training [4].
@@
Analog AirComp relies on uncoded transmission, making it vulnerable to noise and fading, and incompatible with modern digital infrastructures. To address these limitations, several digital AirComp methods have been developed, such as one-bit broadband digital aggregation (OBDA), majority-vote frequency-shift keying (FSK), and their asynchronous and non-coherent extensions. However, these methods are often limited to fewer number of transmitters or specific functions [3]. To provide a more general solution, ChannelComp [5] was introduced to enable arbitrary function computation over a MAC by designing digital modulation schemes via optimization, and it adopts a single-symbol modulation design where each input is mapped to a single modulated symbol.
While ChannelComp is compatible with existing digital communication systems, its single-symbol design limits computation accuracy under noisy conditions and its semidefinite relaxation (SDR) solution involves high optimization complexity. As a result, recent works have focused on multi-symbol modulation designs, which provide additional redundancy and improve robustness [6, 7, 8]. For instance, repetition for multiple access computing (ReMAC[6] enhances computation robustness by selectively repeating modulated symbols across multiple time slots, but this repetition coding design further increases computational complexity. Bit-Slicing [8] offers a lower computational complexity design by partitioning bit sequences into slices and mapping each sliced integer to square quadrature amplitude modulation (QAM) symbols for transmission, yet it is mainly effective for the sum function and suffers from approximation errors when extended to nonlinear computations.
Motivated by these limitations, we propose SeMAC, a new multi-symbol modulation framework that encodes each input into a sequence of symbols using distinct constellation diagrams. As opposed to ReMAC which uses an identical constellation pattern across all time slots, SeMAC leverages constellation diversity to achieve higher computation accuracy. The resulting modulation design problem is formulated as a non-convex quadratically constrained quadratic programming (QCQP), which is NP-hard and computationally demanding [9]. To address this complexity challenge, we develop an iterative algorithm based on SCA combined with SSD, which achieves lower complexity than the standard SDR method. In addition, for scenarios where the modulation format is fixed, we propose SeMAC-PA, which adapts the transmit amplitude and phase of the existing modulation patterns, thereby extending the applicability of our approach to hardware-constrained systems. Numerical results show that SeMAC enables robust computation over a noisy MAC and reduces the computation error by up to 14 dB compared to ReMAC, particularly for the product function.
@@
Figure 1: The overall process of proposed SeMAC via multi-symbol modulation design.
@@ II-A Signal Model
Consider a network with $K$ single-antenna nodes transmitting data to a computation point (CP) over a shared channel. Each node $k$ sends a data value $x_{k}\in\mathbb{F}_{Q}$, where $\mathbb{F}_{Q}$ denotes a finite field with $Q=2^{b}$ elements, where $x_{k}$ is represented using $b$ bits. The goal of the CP is to compute any finite-valued function $f(x_{1},\ldots,x_{K})$ over the MACs. Each node encodes its data value $x_{k}$ into a modulated signal $\vec{x}_{k}$ through an encoder $\mathcal{E}_{k}$, such that $\vec{x}_{k}=\mathcal{E}_{k}(x_{k})\in\mathbb{C}$. We assume perfect synchronization of carrier frequency and symbol timing among all the nodes and the CP [^1] In practical systems, techniques such as frame timing and carrier frequency offset estimation can be applied for achieving synchronization [10]., and then the CP receives the superimposed signal as
@@
Assuming the availability of perfect Channel State Information (CSI), each node applies optimal power control by inverting the channel [11], given by $p_{k}=h_{k}^{*}/|h_{k}|^{2}$. In this regard, we can compensate for channel distortion and ensure coherent signal alignment at the CP. With channel inversion applied, and ignoring the effect of the noise, the received signal $\vec{r}:=\sum\nolimits_{k=1}^{K}\vec{x}_{k}$ yields a finite constellation diagram. Then, the CP applies a tabular function $\mathcal{T}(\cdot)$ to map each constellation point to its corresponding function output.
@@ II-B Overlap Avoidance
$|\vec{v}^{(i)}-\vec{v}^{(j)}|\geq\epsilon|f^{(i)}-f^{(j)}|^{2},\quad\forall(i,j)\in\mathcal{M},$ (2)
where $\mathcal{M}:=\{(i,j)\in[M]^{2}|i<j\}$ and $\epsilon>0$ is a positive constant. Specifically, $M$ is the cardinality of the function range, and $\tilde{M}$ is the number of index pair in set $\mathcal{M}$.
@@ III-A Multi-Symbol Transmission
Furthermore, we construct the modulation matrix $\bm{X}:=[\bm{X}_{1}^{\mathsf{T}},\dots,\bm{X}_{K}^{\mathsf{T}}]^{\mathsf{T}}\in\mathbb{C}^{N\times L}$, where $N:=QK$, by vertically stacking the modulation matrices of all $K$ nodes. Give the input combination $x_{1}^{(i)},\dots,x_{K}^{(i)}$, we define a binary vector $\bm{a}_{i}\in\{0,1\}^{N}$ which consists of $K$ one-hot blocks of length $Q$, each selecting the encoded modulation symbol at one node. Therefore, the resultant modulation sequence associated with the function output $f^{(i)}$ is given by $\bm{v}^{(i)}=\bm{a}_{i}\bm{X}\in\mathbb{C}^{L}$.
@@ III-B Multi-Symbol Estimation and Decoding
$\|(\bm{a}_{i}-\bm{a}_{j})^{\mathsf{T}}\bm{X}\|_{2}^{2}\geq\epsilon|f^{(i)}-f^{(j)}|,\quad\forall(i,j)\in\mathcal{M},$ (6)
where $\|\cdot\|_{2}$ denotes the $\ell_{2}$-norm. Note that if the condition in (6) is not satisfied, the corresponding Voronio regions are merged, and we consider the average of the associated outputs as the final function output [5]. The overall SeMAC process is shown in Figure 1.
@@ IV-A Modulation Design With Channel Inversion
In this subsection, the modulation scheme is designed under the assumption of channel inversion power control. To avoid constellation overlaps as in (6), we pose the following feasibility problem to enable robust computation, subject to the norm power constraint:
$\mathcal{P}_{0}:\quad$ ${\rm find}\quad\bm{X}$
${\rm s.t.}\quad$ ${\rm Tr}(\bm{X}^{\mathsf{H}}\bm{D}_{ij}\bm{X})\geq\Delta f_{ij},\quad\forall(i,j)\in\mathcal{M},$ (7a)
$\|\bm{x}_{k,\ell}\|_{2}^{2}\leq 1,\quad\forall k\in[K],\quad\forall\ell\in[L],$ (7b)
where $\Delta f_{ij}:=\epsilon|f^{(i)}-f^{(j)}|$ and $\bm{D}_{ij}:=(\bm{a}_{i}-\bm{a}_{j})(\bm{a}_{i}-\bm{a}_{j})^{\mathsf{T}}$. However, $\mathcal{P}_{0}$ is a non-convex QCQP, and solving it is NP-hard [5]. A common approach to tackle such problems is the SDR [6], which uses matrix lifting coupled with rank relaxation to relax the problem into a convex semidefinite programming (SDP). While effective for small instances, SDR consumes potentially high complexity for large $K$ and $Q$. To overcome this, we consider a scalable alternative based on SCA to approximate the original problem via a sequence of convex subproblems [12]. Specifically, we first formulate the feasibility problem into an unconstrained minimization problem by hinge-loss penalites:
$\mathcal{P}_{1}:\quad\min_{\bm{X}\in\mathcal{X}}\quad\frac{1}{\tilde{M}}\sum\nolimits_{(i,j)\in\mathcal{M}}g_{ij}(\bm{X}),$ (8)
where $g_{ij}(\bm{X}):=\left[\Delta f_{ij}-\operatorname{Tr}(\bm{X}^{\mathsf{H}}\bm{D}_{ij}\bm{X})\right]^{+}$ penalizes constraint violations and $\mathcal{X}:=\{\|\bm{x}_{k,\ell}\|_{2}^{2}\leq 1,\forall k\in[K],\forall\ell\in[L]\}$ is a convex set of power-constrained vectors. Since $g_{ij}(\bm{X})$ is non-convex, we then approximate it by a convex majorizer at a reference point $\tilde{\bm{X}}^{(s)}$, leading to the following function:
$h_{ij}^{(s)}(\bm{X}):=$ $\left[\Delta f_{ij}-2\Re\{\operatorname{Tr}[(\tilde{\bm{X}}^{(s)})^{\mathsf{H}}\bm{D}_{ij}\bm{X}]\}\right$.
$\left.+{\rm Tr}[(\tilde{\bm{X}}^{(s)})^{\mathsf{H}}\bm{D}_{ij}\tilde{\bm{X}}^{(s)}]\right]^{+}.$ (9)
Hence, at SCA iteration $s$, we solve the following convex subproblem $\mathcal{P}_{2}$, whose solution $\bm{\tilde{X}}^{(s+1)}$ serves as the reference point for the next iteration:
$\mathcal{P}_{2}:\bm{\tilde{X}}^{(s+1)}\in\arg\min_{\bm{X}\in\mathcal{X}}\frac{1}{\tilde{M}}\sum\nolimits_{(i,j)\in\mathcal{M}}h_{ij}^{(s)}(\bm{X}).$ (10)
Algorithm 1 SCA-SSD for solving Problem $\mathcal{P}_{0}$
1:Init: SCA iterations $S$, subgradient steps $T$, step sizes $\{\alpha_{t}\}_{t=0}^{T-1}$. Randomly generate an initial point $\bm{\tilde{X}}^{(0)}\in\mathcal{X}$.
2:for $s=0,1,\dots,S$ do
3:for $t=0,1,\dots,T-1$ do
4:  Sample a single pair $(i_{t},j_{t})$ uniformly from $\mathcal{M}$
5:  $\bm{G}^{(t)}=\partial h^{(s)}_{i_{t}j_{t}}\!\left(\bm{X}^{(t)}\right)$
6:  $\bm{Y}^{(t+1)}=\bm{X}^{(t)}-\alpha_{t}\,\bm{G}^{(t)}$
7:  $\bm{X}^{(t+1)}=\Pi_{\mathcal{X}}\!\left(\bm{Y}^{(t+1)}\right)$
8:end for
9: Set $\bm{\tilde{X}}^{(s+1)}=\bm{X}^{(T)}$
10:end for
11:Output: $\hat{\bm{X}}\leftarrow\bm{\tilde{X}}^{(S+1)}$
Instead of solving $\mathcal{P}_{2}$ by interior-point methods, which consumes high computation cost, we can determine an approximate solution by using the SSD approach. Specifically, at each step $t$, we randomly draw the index pair $(i_{t},j_{t})$ from set $\mathcal{M}$ and apply the following update rule:
$\bm{Y}^{(t+1)}$ $=\bm{X}^{(t)}-\alpha_{t}\cdot\partial h_{i_{t}j_{t}}^{(s)}(\bm{X}^{(t)}),$ (11a)
$\bm{X}^{(t+1)}$ $=\Pi_{\mathcal{X}}(\bm{Y}^{(t+1)}).$ (11b)
where $\alpha_{t}$ is the step size and the subgradient $\partial h_{ij}^{(s)}(\bm{X}^{(t)})$ at the point $\bm{X}^{(t)}$ is computed as:
$\partial h_{ij}^{(s)}(\bm{X}^{(t)})=\begin{cases}-2\bm{D}_{ij}\bm{\tilde{X}}^{(s)},&\text{if }h_{ij}(\bm{X}^{(t)})>0\\ \bm{0},&\textrm{otherwise}.\end{cases}$ (12)
Moreover, $\Pi_{\mathcal{X}}(\cdot)$ denotes the projection onto $\mathcal{X}$ given by:
$\bm{x}_{k,\ell}:=\begin{cases}\bm{y}_{k,\ell},&\text{if }\|\bm{y}_{k,\ell}\|_{2}\leq 1,\\ \frac{\bm{y}_{k,\ell}}{\|\bm{y}_{k,\ell}\|_{2}},&\text{otherwise}.\end{cases}$ (13)
The complete process is summarized in Algorithm 1, which integrates the SCA with SSD updates.
@@
Note that solving the standard SDR problems require $\mathcal{O}(\tilde{M}^{4}N^{0.5}L^{0.5})$ operations due to matrix lifting [6], which becomes computationally demanding for large $K$ or $Q$. In contrast, the proposed SCA–SSD algorithm samples only one constraint per update, leading to per-iteration complexity $\mathcal{O}(NL)$ and total cost $\mathcal{O}(STNL)$. This linear scaling avoids dependence on $\tilde{M}$ and thereby ensures scalability in larger networks. Furthermore, with a diminishing step-size schedule, the SSD updates follow the expected descent direction and the framework converges in expectation to a stationary point [13].